OpenStudy (anonymous):

Reduce 7x exponent 2 divided by 14x

OpenStudy (anonymous):

\[\frac{ 7x^2 }{ 14x }\]?

OpenStudy (studygurl14):

\(\Large\frac{7x^2}{14x}\)

OpenStudy (studygurl14):

\(\large \frac{7x^2}{14x}\rightarrow\frac{7x(x)}{7x(2)}\rightarrow\frac{x}{2}\) (the 7x's cancel each other out)

OpenStudy (studygurl14):

Does that make sense @SilverPersia